The temperature of an ideal gas remains constant while the absolute pressure changes from 100 kpa to 800 kpa. If initial volume is 100 liters, what is the final volume in liters?



1
Expert's answer
2022-03-24T16:38:24-0400

Ideal gas law

P1V1 = nRT1 ________eqn(1)

P2V2 = nRT2 _________eqn(2)

From the given conditions,

Pressures and Volumes are vary, but the Temperatures, number of moles remains constant.

And R is the universal gas constant

So,the above equations becomes

"P_{1}V_{1}=P_{2}V_{2}\>\>equation (3)"

"P_{1}=100\>kpa\\P_{2}=800\>kpa\\V_{1}=100\>liters\\V_{2}=final\>Volume=?"

Put all the above values in equation (3)

"P_{1}V_{1}=P_{2}V_{2}"

"{100kpa}\times{100liters}={800kpa}\times{V_{2}}"

"V_{2} = [\frac{100kpa}{800kpa}][100liters]\\V_{2} = 12.5\>liters"

"V_{2}=12.5\>liters"

So, the final Volume V2 is 12.5 liters